# CMB Constraints on Loop-Induced Decays of Leptophilic Dark Matter

### v1

@@ II Models
We remain agnostic about the specific DM production mechanism. Instead, we focus on the low-energy phenomenology of DM at later times, around the CMB era, allowing our study to be effectively independent of the UV completion of the DM models. However, we do assume DM is produced non-thermally, and DM must remain thermally decoupled from the SM before and around the time of BBN to avoid constraints on light element abundances from BBN and CMB anisotropies [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. For each model of interest, we verify that our upper bounds on the strength of DM-electron couplings are sufficiently small such that DM remains non-thermalized with the SM bath at temperatures $T\gtrsim\mathrm{MeV}$.
@@ II.1 Scalar
where $g_{\phi\bar{e}e}$ is the scalar DM-electron coupling constant. At leading order, $\phi$ decays into two photons through an electron loop, with corresponding decay width
@@ III Energy injection and deposition
Figure 1: Residuals in the hydrogen ionization channel of the energy deposition function, $f_{\mathrm{H\,ion}}(z)$, for $m_{\mathrm{DM}}=10\,\mathrm{keV}$ and a given cosmology relative to the Planck 2018 best-fit baseline. The three panels (top to bottom) correspond to scalar DM-electron couplings $g_{\phi\bar{e}e}$ at the 95% C.L. exclusion limit from Ref. [40] (top), an order of magnitude larger (middle), and an order of magnitude smaller (bottom). In each panel, the eight curves, shown in various shades of red, represent different combinations of $(h,\,\omega_{b},\,\omega_{c})$ taken at their ±$\pm 95\%$ C.L. Planck 2018 values.
@@
In Fig. 1, we show how these cosmological parameter variations affect the hydrogen ionization channel, which is the most relevant in setting CMB constraints on decaying DM, in the energy deposition function. We plot residuals in $f_{\mathrm{H_{ion}}}(z)$ for eight combinations of $(h,\,\omega_{b},\,\omega_{c})$, each taken at their ±$\pm 95\%$ C.L. Planck 2018 limits, relative to the fiducial (Planck 2018 best-fit) cosmology used in DarkHistory. As a representative example, we take $m_{\mathrm{DM}}=10\,\mathrm{keV}$ and consider three choices of scalar DM-electron coupling $g_{\phi\bar{e}e}$: the first yields a DM lifetime at the 95% C.L. exclusion limit reported in Ref. [40], while the other two are an order of magnitude larger and smaller than this baseline value. Equivalent plots for the other DM masses relevant to our analysis are provided for completeness in Appendix A. Overall, we observe $\sim\mathcal{O}(1)\%$ residual changes in $f_{\rm H_{ion}}(z)$, which remain roughly independent of $g_{\phi\bar{e}e}$ near recombination but grow slightly larger at lower redshifts for stronger couplings. These results support the expectation that cosmology-driven variations in $f_{c}(z)$ are modest [41], yet they motivate a dedicated check to ensure that such percent-level corrections to the $f_{c}(z)$ do not bias the final DM-electron coupling limits derived from a full MCMC treatment. This verification has not been explicitly performed in previous studies, and we address it in the next section.
@@ IV Analysis
Regarding aspect (ii), the recent NPIPE maps offer advancements compared to previous Planck data releases, including increased sky coverage at high frequencies and improved processing of time-ordered data, which lead to a reduction in small-scale noise. These improvements provide 10% to 20% smaller uncertainties on constraints of $\Lambda$CDM parameters and minimal extensions [49], further motivating their application to decaying DM models.
@@
We use the latest NPIPE Planck data release and use the corresponding Planck PR4 likelihood code [45]. We replace the Planck PR3 plik-lite $-\ell$ likelihood with the updated HiLLiPoP likelihood [49]. For the polarization spectra at $\ell<30$ we use the PR4 LoLLiPoP likelihood [49] instead of the PR3 SimAll likelihood, while retaining the $-\ell$ Commander likelihood for temperature power spectra and the SMICA lensing reconstruction likelihood. [^2] An updated PR4 lensing likelihood based on the NPIPE data release is also available [51], but we do not include it in our analysis. Decaying DM has negligible impact on the lensing convergence signal when the decay time is significantly longer than the age of the universe [52]. Incorporating this likelihood would only marginally tighten constraints on the DM density, with no appreciable effects on our exclusions limits. As described for PR3-$f^{\mathrm{FC}}_{c}$, the $f_{c}(z)$ functions are precomputed by DarkHistory under a fixed fiducial cosmology.
Figure 2: Planck 95% C.L. upper limits on the dimensionless DM-electron coupling in the scalar ($g_{\phi\bar{e}e}$) (left panel, left axis), pseudo-scalar ($g_{\Phi\bar{e}e}$) (left panel, right axis), and ALP ($g_{a\bar{e}e}$) (right panel) cases as a function of the DM mass, $m_{\rm{DM}}$. For each panel, the black curve shows the results from the Planck 2018 PR3 likelihood with the energy deposition functions $f_{c}(z)$ from Eq. (8) pre-computed under a fiducial $\Lambda$CDM cosmology (PR3-$f_{c}^{\rm{FC}}$); the blue curve corresponds to the PR3 analysis for which the $f_{c}(z)$ functions are evaluated at each MCMC step to reflect the sampled cosmological parameters (PR3-$f_{c}^{\rm{CC}}$); and the red curve corresponds to the constraints obtained with the updated Planck PR4 likelihood using the precomputed $f_{c}(z)$ functions from a fiducial cosmology (PR4-$f_{c}^{\rm{FC}}$).
@@ V Results
As expected, our results from the standard, fixed-cosmology PR3-$f_{c}^{\rm{FC}}$ analysis are consistent with the bounds on DM decay reported in previous studies [39, 38, 40]. Importantly, our results for the cosmology-consistent PR3-$f_{c}^{\rm{CC}}$ analysis aligns with those from fixed-cosmology approach at the 10% level. This agreement validates the standard methodology of assuming a fixed fiducial $\Lambda$CDM cosmology when pre-computing the $f_{c}(z)$ functions from DarkHistory for decaying DM. It further highlights that percent-level differences in the computation of the energy deposition functions do not have any appreciable impact on our DM decay bounds. As discussed in Appendix B, we observe that the cosmology-consistent approach reduces mild parameter degeneracies present in the fixed-cosmology case, particularly at low DM masses, though this reduction has negligible impact on the derived constraints.
We find that the upper limits derived from the Planck PR4 analysis are consistent with those obtained using the PR3 likelihoods, with only minor differences observed for $m_{\rm{DM}}\geq 10\,$keV. These $20\%$ improvements in the DM-electron coupling constant exclusion limits can be attributed to lower noise levels in the LoLLiPoP likelihood used for $-\ell$ polarization spectra in PR4. However, for $m_{\rm{DM}}\lesssim 10\,$keV, the constraints are effectively the same, likely due to increased degeneracy between $g_{\phi\bar{e}e}$, $\ln(10^{10}A_{s})$, and $z_{\rm{reio}}$ in this mass range, which further limits the impact of the improved polarization data (see also Appendix B). In general, the lack of significant improvements in the upper limits reflects the fact that most of the constraining power for decaying DM originates from $-\ell$ polarization data, where the differences between PR3 and PR4 are modest, limiting the potential gains from the enhanced $-\ell$ precision in the NPIPE data release.
@@ V.1 Flavorful Variations
Our framework also applies to the cases of 1-loop muon- and tau-mediated decays; therefore, models in which DM couples exclusively to muons or taus are also constrained by our analysis. Such DM models are well-motivated  [57, 58, 59] and open up a host of collider [60, 61] and astrophysical probes [62, 63, 48]. The fact that such models are constrained by the CMB even though they do not couple to electrons (i.e., no additional tree-level scatterings processes are present at all during the epoch of recombination) further highlights the importance of the 1-loop decay. Decay widths in these scenario are obtained with a simple rescaling $m_{e}\rightarrow m_{\mu}$ or $m_{\tau}$ in Eqs. (2), (4) and (6). In other words, the bounds presented in Fig. 2 for the DM-electron coupling equally apply to the DM-muon or DM-tau coupling scenarios, after rescaling them by a factor $m_{\mu}/m_{e}\sim 10^{3}$ and $m_{\tau}/m_{e}\sim 10^{5}$, respectively.
@@ VI Comparison with other constraints
In Figs. 3, 4, and 5, we show our strongest upper limits on scalar, pure pseudo-scalar, and ALP DM-electron couplings from our Planck PR4 analysis and compare them with existing bounds in the literature.
For the scalar case shown in Fig. 3, stellar cooling constraints from horizontal branch (HB) and red giants (RG) stars [64] provided the most stringent bounds on $g_{\phi\bar{e}e}$ for $m_{\phi}\lesssim 50\,$keV prior to our work. For masses larger than $50$ keV, we re-cast constraints on the non-detection of X-ray and gamma-ray signals from the NuStar, XMM-Newton, and INTEGRAL telescopes [65, 66, 67, 68] produced by $\phi\to\gamma\gamma$ decays. Exclusion limits from CMB anisotropies derived in this work, shown in blue, surpass previous bounds for $1\,\text{keV}\lesssim m_{\phi}\lesssim 5\,\text{keV}$ by up to an order of magnitude in this range. Although our CMB bounds are weaker than the astrophysical bounds in most of the mass range shown, they are robust and are not subject to the same astrophysical uncertainties.
@@
In the case of pseudo-scalars and ALPs shown in Figs. 4 and 5, respectively, DM-electron interactions are equivalent at tree-level. Consequently, terrestrial limits and stellar constraints are identical for the two classes of models, although we display only a subset of these bounds in each figure to highlight the parameter space most relevant for our derived CMB constraints. Specifically, for the pure pseudo-scalar case in Fig. 4, we show the direct detection bounds from DarkSide [70], PandaX [71, 72] and XENON [73, 74, 75, 76], along with the stellar cooling constraints from RG stars [64, 69]. For ALPs in Fig. 5, we also include the SuperCDMS [77], EDELWEISS [78], and GERDA [79] exclusion limits, as well as the solar luminosity/neutrino bound [63].
@@ VII Conclusions
We find that the cosmology-consistent analysis reproduces the results obtained using the standard fixed-cosmology approach, validating the latter methodology for deriving constraints on DM decay. This outcome is important for future studies, since running a cosmology-consistent analysis with DarkHistory takes significantly more computational time and resources than running a fixed-cosmology analysis. Our results using the Planck PR3 and PR4 likelihoods are largely consistent, with only minor improvements observed for $m_{\rm{DM}}\gtrsim 10\,$keV. This consistency reflects the fact that constraints on decaying DM are predominantly driven by $-\ell$ polarization data, where the improvements introduced by the PR4 likelihoods are modest.
@@
VIII Acknowledgments
We are grateful to Hongwan Liu and Wenzer Qin for useful discussions. We acknowledge the use of CLASS [44], GetDist [56], IPython [80], MontePython [53, 54], and the Python packages Matplotlib [81], NumPy [82] and SciPy [83]. We also acknowledge the use of the AxionLimits Github repository [84],which was an invaluable resource in verifying existing axion constraints and extracting most of the bounds from the current literature shown in Figs. 4 and 5.
NB is supported by PRD/ARPE 2022 “Cosmology with Gravitational waves and Large Scale Structure - CosmoGraLSS”. KB acknowledges support from the National Science Foundation under Grant No. PHY-2413016. GM acknowledges support by the U.S. Department of Energy, Office of Science, Office of High Energy Physics program under Award Number DE-SC-0022021, as well as the Swedish Research Council (Contract No. 638-2013-8993). GM is also supported by the Continuing Fellowship of the Graduate School of the College of Natural Sciences at the University of Texas at Austin.
@@ Appendix A Cosmology Dependence of the energy deposition functions
Figure 7: Marginalized 2D posterior distributions for the $\Lambda$CDM parameters and the dimensionless DM-electron coupling, $g_{\phi\bar{e}e}$, for a DM mass of 111\,1keV. Contours correspond to the 68% and 95% confidence levels, derived from a joint analysis of the Planck PR4 likelihoods, incorporating temperature, polarization, and lensing anisotropies, as described in Sec. IV. The 1D marginalized posteriors for each parameter are displayed along the diagonal.
@@ Appendix B Posterior Distributions and Parameter Degeneracies
In this appendix, we present representative examples of the reconstructed marginalized posterior probability distributions from our MCMC analyses and comment on degeneracies between parameters. We show in Fig. 7 the triangle plot of the 2D posterior distributions for the $\Lambda$CDM parameters {$10^{2}\omega_{b}$, $\omega_{c}$, $10^{2}\theta_{s}$, $\ln(10^{10}A_{s})$, $n_{s}$ and $z_{\rm reio}$} and the dimensionless DM-electron coupling $g_{\phi\bar{e}e}$, for a DM mass of $1\,\rm{keV}$. These results are derived from the Planck PR4 likelihood analysis (PR4-$f^{\mathrm{FC}}_{c}$), described in Sec. IV, where we follow the standard treatment of energy injection, using pre-computed energy deposition functions $f_{c}(z)$ under a fiducial cosmology.
As expected, the constraints on the standard $\Lambda$CDM parameters remain consistent with those from a $\Lambda$CDM-only analysis of the Planck PR4 data [49]. In addition, we note a mild degeneracy between $z_{\rm reio}$, $g_{\phi\bar{e}e}$, and $\ln(10^{10}A_{s})$. The end of reionization is delayed in the presence of additional processes that contribute to ionizing gas, leading to a negative correlation between $g_{\phi\bar{e}e}$ and $z_{\rm reio}$. Although one might expect extra ionization to anticipate the end of reionization, partial ionization from DM decays at earlier times effectively reduces the astrophysical reionization required at high redshift. In the MCMC fit, this translates into a lower $z_{\rm reio}$ to avoid overestimating the total optical depth, leading to a small negative correlation between $g_{\phi\bar{e}e}$ and $z_{\rm reio}$. Furthermore, the optical depth at reionization is closely tied to the overall amplitude of the CMB anisotropies, creating the known positive correlation between the amplitude of the primordial scalar power spectrum $A_{s}$ and $z_{\rm reio}$. As a result, a weak degeneracy between $g_{\phi\bar{e}e}$ and $\ln(10^{10}A_{s})$ is induced through their mutual dependence on $z_{\rm reio}$.
In Fig. 8 (left panel), we further explore these degeneracies by comparing the joint posterior distributions of $z_{\rm reio}$, $g_{\phi\bar{e}e}$, and $\ln(10^{10}A_{s})$ within the same PR4-$f^{\rm FC}_{c}$ analysis for DM masses of 111\,1keV, 101010\,10keV, and 100100100\,100keV. We rescale DM-electron coupling for each mass by the respective 95% C.L. upper limit [i.e., we show $\log_{10}(g_{\phi\bar{e}e}/g_{\phi\bar{e}e}^{95\%\,\rm{C.L.}})$], to more easily compare the 1D and 2D posterior distributions in a single plot. Overall, this comparison reveals that the degeneracies among $z_{\rm reio}$$g_{\phi\bar{e}e}$, and $\ln(10^{10}A_{s})$ gradually diminish as the DM mass increases, manifesting as flatter 1D posteriors for $g_{\phi\bar{e}e}$ at values well below the 95% C.L. upper limit for $m_{\rm{DM}}\gtrsim 10\,\rm{keV}$.

### v2

@@ II Models
We remain agnostic about the specific DM production mechanism. Consequently, we treat the relic abundance as having been set earlier by an independent non-thermal process (e.g., field misalignment), rather than by the electron coupling explored here. Instead, we focus on the low-energy phenomenology of DM at later times, around the CMB era, allowing our study to be effectively independent of the UV completion of the DM models. However, we do assume DM is produced non-thermally, and DM must remain thermally decoupled from the SM before and around the time of BBN to avoid constraints on light element abundances from BBN and CMB anisotropies [16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. For each model of interest, we verify that our upper bounds on the strength of DM-electron couplings are sufficiently small such that DM remains non-thermalized with the SM bath at temperatures $T\gtrsim\mathrm{MeV}$.
@@
$\phi,\Phi,a$$\gamma$$\gamma$
Figure 1: Feynman diagram for the electron-mediated one-loop decay of scalar ($\phi$), pseudo-scalar ($\Phi$), or axion-like particle ($a$) DM into two photons. This decay channel represents the dominant interaction responsible for observable cosmological signatures analyzed in this work.
@@ II.1 Scalar
where $g_{\phi\bar{e}e}$ is the scalar DM-electron coupling constant. At leading order, $\phi$ decays into two photons through an electron loop, illustrated in Fig. 1, with corresponding decay width
@@ III Energy injection and deposition
Figure 2: Residuals in the hydrogen ionization channel of the energy deposition function, $f_{\mathrm{H\,ion}}(z)$, for $m_{\mathrm{DM}}=10\,\mathrm{keV}$ and a given cosmology relative to the Planck 2018 best-fit baseline. The three panels (top to bottom) correspond to scalar DM-electron couplings $g_{\phi\bar{e}e}$ at the 95% C.L. exclusion limit from Ref. [40] (top), an order of magnitude larger (middle), and an order of magnitude smaller (bottom). In each panel, the eight curves, shown in various shades of red, represent different combinations of $(h,\,\omega_{b},\,\omega_{c})$ taken at their $\pm 95\%$ C.L. Planck 2018 values.
@@
In Fig. 2, we show how these cosmological parameter variations affect the hydrogen ionization channel, which is the most relevant in setting CMB constraints on decaying DM, in the energy deposition function. We plot residuals in $f_{\mathrm{H_{ion}}}(z)$ for eight combinations of $(h,\,\omega_{b},\,\omega_{c})$, each taken at their $\pm 95\%$ C.L. Planck 2018 limits, relative to the fiducial (Planck 2018 best-fit) cosmology used in DarkHistory. As a representative example, we take $m_{\mathrm{DM}}=10\,\mathrm{keV}$ and consider three choices of scalar DM-electron coupling $g_{\phi\bar{e}e}$: the first yields a DM lifetime at the 95% C.L. exclusion limit reported in Ref. [40], while the other two are an order of magnitude larger and smaller than this baseline value. Equivalent plots for the other DM masses relevant to our analysis are provided for completeness in Appendix A. Overall, we observe $\sim\mathcal{O}(1)\%$ residual changes in $f_{\rm H_{ion}}(z)$, which remain roughly independent of $g_{\phi\bar{e}e}$ near recombination but grow slightly larger at lower redshifts for stronger couplings. These results support the expectation that cosmology-driven variations in $f_{c}(z)$ are modest [41], yet they motivate a dedicated check to ensure that such percent-level corrections to the $f_{c}(z)$ do not bias the final DM-electron coupling limits derived from a full MCMC treatment. This verification has not been explicitly performed in previous studies, and we address it in the next section.
@@ IV Analysis
[…]
@@
We use the latest NPIPE Planck data release and use the corresponding Planck PR4 likelihood code [45]. We replace the Planck PR3 plik-lite high-$\ell$ likelihood with the updated HiLLiPoP likelihood [49]. For the polarization spectra at $\ell<30$ we use the PR4 LoLLiPoP likelihood [49] instead of the PR3 SimAll likelihood, while retaining the low-$\ell$ Commander likelihood for temperature power spectra and the SMICA lensing reconstruction likelihood. [^2] An updated PR4 lensing likelihood based on the NPIPE data release is also available [51], but we do not include it in our analysis. Decaying DM has negligible impact on the lensing convergence signal when the decay time is significantly longer than the age of the universe [52]. Incorporating this likelihood would only marginally tighten constraints on the DM density, with no appreciable effects on our exclusions limits. As described for PR3-$f^{\mathrm{FC}}_{c}$, the $f_{c}(z)$ functions are precomputed by DarkHistory under a fixed fiducial cosmology.
@@
For each dataset, we sample the following cosmological parameters with broad flat priors: the angular size of the sound horizon at decoupling $\theta_{s}$, the physical baryon density $\omega_{b}$, the physical matter density $\omega_{m}$, the scalar amplitude $\ln(10^{10}A_{s})$, the scalar spectral index $n_{s}$, and the redshift at reionization $z_{\rm{reio}}$. Additionally, we sample $\log_{10}\left(g_{\phi\bar{e}e}\right)$ with a flat prior for fixed DM masses $m_{\mathrm{DM}}=\{100\,\rm{eV},\,1\,\rm{keV},\,10\,\rm{keV},\,100\,\rm{keV},\,1\,\rm{MeV}\}$. Since the structures of the pseudo-scalar and ALP decay widths are formally equivalent to that of scalar DM, it is possible to recast scalar DM upper limits into pseudo-scalar and ALP limits via the rescalings $g_{\phi\bar{e}e}\rightarrow\sqrt{288/64}\,g_{\Phi\bar{e}e}$ and $g_{\phi\bar{e}e}\rightarrow\sqrt{288/64}\,m_{\rm{DM}}^{2}/(12\,m_{e}^{2})\,g_{a\bar{e}e}$, respectively.
@@ V Results
[…]
We find that the upper limits derived from the Planck PR4 analysis are consistent with those obtained using the PR3 likelihoods, with only minor differences observed for $m_{\rm{DM}}\geq 10\,$keV. These $20\%$ improvements in the DM-electron coupling constant exclusion limits can be attributed to lower noise levels in the LoLLiPoP likelihood used for low-$\ell$ polarization spectra in PR4. However, for $m_{\rm{DM}}\lesssim 10\,$keV, the constraints are effectively the same, likely due to increased degeneracy between $g_{\phi\bar{e}e}$, $\ln(10^{10}A_{s})$, and $z_{\rm{reio}}$ in this mass range, which further limits the impact of the improved polarization data (see also Appendix B). In general, the lack of significant improvements in the upper limits reflects the fact that most of the constraining power for decaying DM originates from low-$\ell$ polarization data, where the differences between PR3 and PR4 are modest, limiting the potential gains from the enhanced high-$\ell$ precision in the NPIPE data release.
@@ V.1 Flavorful Variations
Our framework also applies to the cases of 1-loop muon- and tau-mediated decays; therefore, models in which DM couples exclusively to muons or taus are also constrained by our analysis. Such DM models are well-motivated  [57, 58, 59] and open up a host of collider [60, 61] and astrophysical probes [62, 63, 48, 64]. The fact that such models are constrained by the CMB even though they do not couple to electrons (i.e., no additional tree-level scatterings processes are present at all during the epoch of recombination) further highlights the importance of the 1-loop decay. Decay widths in these scenario are obtained with a simple rescaling $m_{e}\rightarrow m_{\mu}$ or $m_{\tau}$ in Eqs. (2), (4) and (6). In other words, the bounds presented in Fig. 3 for the DM-electron coupling equally apply to the DM-muon or DM-tau coupling scenarios, after rescaling them by a factor $m_{\mu}/m_{e}\sim 10^{3}$ and $m_{\tau}/m_{e}\sim 10^{5}$, respectively.
@@ VI Comparison with other constraints
In Figs. 4, 5, and 6, we show our strongest upper limits on scalar, pure pseudo-scalar, and ALP DM-electron couplings from our Planck PR4 analysis and compare them with existing bounds in the literature.
For the scalar case shown in Fig. 4, stellar cooling constraints from horizontal branch (HB) and red giants (RG) stars [65, 70] provided the most stringent bounds on $g_{\phi\bar{e}e}$ for $m_{\phi}\lesssim 50\,$keV prior to our work. For masses larger than $50$ keV, we re-cast constraints on the non-detection of X-ray and gamma-ray signals from the NuStar, XMM-Newton, and INTEGRAL telescopes [66, 67, 68, 69] produced by $\phi\to\gamma\gamma$ decays. Exclusion limits from CMB anisotropies derived in this work, shown in blue, surpass previous bounds for $1\,\text{keV}\lesssim m_{\phi}\lesssim 5\,\text{keV}$ by up to an order of magnitude in this range. Although our CMB bounds are weaker than the astrophysical bounds in most of the mass range shown, they are robust and are not subject to the same astrophysical uncertainties.
@@
In the case of pseudo-scalars and ALPs shown in Figs. 5 and 6, respectively, DM-electron interactions are equivalent at tree-level. Consequently, terrestrial limits and stellar constraints are identical for the two classes of models, although we display only a subset of these bounds in each figure to highlight the parameter space most relevant for our derived CMB constraints. Specifically, for the pure pseudo-scalar case in Fig. 5, we show the direct detection bounds from DarkSide [72], PandaX [73, 74] and XENON [75, 76, 77, 78], along with the stellar cooling constraints from RG stars [65, 71]. For ALPs in Fig. 6, we also include the SuperCDMS [79], EDELWEISS [80], and GERDA [81] exclusion limits, as well as the solar luminosity/neutrino bound [63].
@@ VII Conclusions
We find that the cosmology-consistent analysis reproduces the results obtained using the standard fixed-cosmology approach, validating the latter methodology for deriving constraints on DM decay. This outcome is important for future studies, since running a cosmology-consistent analysis with DarkHistory takes significantly more computational time and resources than running a fixed-cosmology analysis. Our results using the Planck PR3 and PR4 likelihoods are largely consistent, with only minor improvements observed for $m_{\rm{DM}}\gtrsim 10\,$keV. This consistency reflects the fact that constraints on decaying DM are predominantly driven by low-$\ell$ polarization data, where the improvements introduced by the PR4 likelihoods are modest.
@@
Acknowledgements.
We are grateful to Can Kilic, Hongwan Liu and Wenzer Qin for useful discussions. We acknowledge the use of CLASS [44], GetDist [56], IPython [82], MontePython [53, 54], and the Python packages Matplotlib [83], NumPy [84] and SciPy [85]. We also acknowledge the use of the AxionLimits Github repository [86],which was an invaluable resource in verifying existing axion constraints and extracting most of the bounds from the current literature shown in Figs. 5 and 6. We acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing high-performance computing resources that have contributed to the research results reported within this paper. NB is supported by PRD/ARPE 2022 “Cosmology with Gravitational waves and Large Scale Structure - CosmoGraLSS”. KB acknowledges support from the National Science Foundation under Grant No. PHY-2413016. GM acknowledges support by the U.S. Department of Energy, Office of Science, Office of High Energy Physics program under Award Number DE-SC-0022021, as well as the Swedish Research Council (Contract No. 638-2013-8993). GM is also supported by the Continuing Fellowship of the Graduate School of the College of Natural Sciences at the University of Texas at Austin.
@@ Appendix A Cosmology Dependence of the energy deposition functions
In this appendix, we present for completeness the cosmology-driven variations in the hydrogen ionization channel of the energy deposition function, $f_{\rm H_{ion}}(z)$, for the four additional DM masses analyzed in this work, namely $m_{\mathrm{DM}}$ of $100\,\mathrm{eV}$, $1\,\mathrm{keV}$, $100\,\mathrm{keV}$, and $1\,\mathrm{MeV}$. In Fig. 7 (analogous to Fig. 2), we display, for each DM mass, the residuals in $f_{\rm H_{ion}}(z)$ computed under eight different cosmologies, with $(h,\,\omega_{b},\,\omega_{c})$ taken at their $\pm 95\%$ C.L. Planck 2018 limits. For each mass, we consider three representative choices of the scalar DM–electron coupling, $g_{\phi\bar{e}e}$: a value that yields a DM lifetime at the 95% C.L. exclusion limit from Ref. [40] and values that are one order of magnitude larger and smaller. While the precise residuals depend on both the DM mass and coupling (e.g., slightly larger deviations for DM masses >1>1\,keV and for stronger couplings at lower redshifts), we consistently find $f_{\rm H_{ion}}(z)$ changes remain at roughly the $\mathcal{O}(1)\%$ level near recombination. These small but non-zero departures motivate our cosmology-consistent analysis, $\textbf{PR3-}f^{\rm CC}_{c}$, although a comprehensive examination of their detailed origins lies beyond the scope of this work. In practice, the outcome of our $\textbf{PR3-}f^{\rm CC}_{c}$ study (see Sec. V) demonstrates that these mild variations have no impact on the derived DM–electron coupling constraints and can thus be safely neglected in subsequent analyses of DM decay.
@@ Appendix B Posterior Distributions and Parameter Degeneracies
[…]
In Fig. 9 (left panel), we further explore these degeneracies by comparing the joint posterior distributions of $z_{\rm reio}$, $g_{\phi\bar{e}e}$, and $\ln(10^{10}A_{s})$ within the same PR4-$f^{\rm FC}_{c}$ analysis for DM masses of 11\,keV, 1010\,keV, and 100100\,keV. We rescale DM-electron coupling for each mass by the respective 95% C.L. upper limit [i.e., we show $\log_{10}(g_{\phi\bar{e}e}/g_{\phi\bar{e}e}^{95\%\,\rm{C.L.}})$], to more easily compare the 1D and 2D posterior distributions in a single plot. Overall, this comparison reveals that the degeneracies among $z_{\rm reio}$$g_{\phi\bar{e}e}$, and $\ln(10^{10}A_{s})$ gradually diminish as the DM mass increases, manifesting as flatter 1D posteriors for $g_{\phi\bar{e}e}$ at values well below the 95% C.L. upper limit for $m_{\rm{DM}}\gtrsim 10\,\rm{keV}$.